OpenStudy (anonymous):

Write an equation of the line with the given slope and y-intercept slope: 7/6, y-intercept: 3

OpenStudy (anonymous):

a.y=7/6x+3 b.y=-7/6x+3 c.y=7/6x-3 d.y=6/7x+3